The number density of electrons in a metal is 28 3 9.64 10 m− × . What is the drift
velocity of electrons in a wire made up of this metal when a current of 4A flows in
it given that its cross-sectional area is 2.0 mm2
? How long will it take for an
electron to travel a distance of 0.10 m in this wire?
1
Expert's answer
2019-04-01T10:39:24-0400

As far as the current I = vSne, then the drift velocity

v = I/Sne = 4/(2*10-6*9.64*1028*1.6*10-19) = 0.13 (mm/s), respectively Ï„ = L/v = 10/0.13 = 77 (s).

So, v = 0.13 mm/s, τ = 77 s.